OpenStudy (anonymous):

Particle A moves along the positive horizontal axis, and particle B along the graph of \(f(x)=-\sqrt{3}x,x\le 0\). At a certain time, A is at the point (5,0) and moving with speed 3 units/sec; and B is at a distance of 3 units from the origin and moving with speed 4 units/sec. At what rate is the distance between A and B changing?

OpenStudy (inkyvoyd):

Nope, too hard.

OpenStudy (anonymous):

Don't be a quitter inky haha

OpenStudy (anonymous):

Are the coordinates from which they start given????

OpenStudy (anonymous):

I posted the question text in full, that's all the information I have.

OpenStudy (kinggeorge):

I think the first thing to do is find what points on the line \(y=-\sqrt3 x\) are a distance of 3 from the origin.

OpenStudy (anonymous):

Well, it's restricted to \(x\le0\) so it's just the one point, haven't solved for the exact point yet.

OpenStudy (anonymous):

\(x^2+y^2=9\) and \(y=-\sqrt{3}x\), so if we solve that system we get... an imaginary answer. But we can actually change that to be \(y=\sqrt{3}x\) and then flip the sign later, and we get x=-3/2, and y is \(3\sqrt{3}/2\), right?

OpenStudy (kinggeorge):

That looks right.

OpenStudy (anonymous):

So we know the exact position but I still have no idea how to figure out a formula for the distance between A and B so as to have something to derive to determine a rate of change.

OpenStudy (kinggeorge):

Now we want to find a parameterization for the lines.

OpenStudy (kinggeorge):

So for particle A:\[\binom{5}{0}+t\binom{3}{0}\]

OpenStudy (kinggeorge):

And for particle B it's more complicated.

OpenStudy (anonymous):

Hrmm this seems almost like a vector problem, finding the y and x components of the vector B

OpenStudy (kinggeorge):

Almost.

OpenStudy (anonymous):

But we haven't done anything like that in this book so I don't think that's how he wants me to do it

OpenStudy (kinggeorge):

Well, let's find the distance between the two point right now.

OpenStudy (anonymous):

Right now the distance between them is 7

OpenStudy (anonymous):

Can we find distance after a small time 'dt' and then find the distance again, and then find the rate of change of distance????

OpenStudy (kinggeorge):

That would be the basic approach.

OpenStudy (anonymous):

I guess we'll have to use that only because we have no other info as to where the particles are starting, do the have uniform velocity.....

OpenStudy (anonymous):

is the answer 83/14 ????

OpenStudy (anonymous):

I don't have the answer, it's not in the book.

OpenStudy (kinggeorge):

We would want to find a relation between the time t and and the distance between the two particles where t=0 corresponds to where the points are located now.

OpenStudy (anonymous):

Yep. I know that's what we need to do I just have no idea how to do it.

OpenStudy (kinggeorge):

You could always just find the distance when A, B are just a wee bit farther along, and estimate from that XP

OpenStudy (anonymous):

Haha I want to learn the right method for this, though. It's in a chapter about Differentiation techniques, sandwiched in between two problems on the Chain Rule. There's gotta be an easier way to do this I'm just missing it.

OpenStudy (kinggeorge):

I'm far too tired to think.

OpenStudy (anonymous):

It's a question of differentials. Consider the situation after time 'dt', the new positions of the particles would be: A : (5 + 3 dt , 0) B : (-1.5 - 2 dt , 3^(3/2)/2 + 2(3^.5) dt) Now, find the distance between them. You can neglect (dt)^2, since dt is very small. The rate of change of distance is change in distance / dt. Now, limit of the ratio as dt->0 should give you the answer, i guess, by the L'Hospitals Rule.

OpenStudy (anonymous):

Alright, I did it by hand, after t they are at positions \((-3.5, 3.5\sqrt{3})\) and (8,0), and the distance between them is 13. So the distance changed by 6. The answer isn't 6, though, is it?

OpenStudy (kinggeorge):

Probably not.

OpenStudy (anonymous):

After time t, the positions should be in terms of t.....

OpenStudy (anonymous):

Okay I actually didn't do after time t, I did after one second. After one second the distance increases by 6.

OpenStudy (anonymous):

I guess we have to find the rate of change of distance at a particular instant, so we should use d/dt (distance between A and B)

OpenStudy (anonymous):

Thats where your differentiation will come in.....

OpenStudy (dumbcow):

 we know the angle is 120 because tan(theta) = sqrt3 use law of cosines to find distance (c) in terms of lengths a and b Note: cos(120) = -1/2 \[c = \sqrt{a^{2}+b^{2}+ab}\] Now the goal is to find rate distance is changing or dc/dt \[\frac{dc}{dt} = \frac{da}{dt}*\frac{dc}{da}+\frac{db}{dt}*\frac{dc}{db}\] da/dt = 3 db/dt = 4 \[\frac{dc}{da}=\frac{2a+b}{2\sqrt{a^{2}+b^{2}+ab}}\] \[\frac{dc}{db}=\frac{a+2b}{2\sqrt{a^{2}+b^{2}+ab}}\] only thing left is plug in values for a,b which are the given lengths a = 5 b = 3 this will give the rate the distance is changing at that moment in time i believe Ashu was correct...83/14 hope this made sense